write an equation and solve.
Dan opened a piggy bank and found a total of thirty-five coins, made up of nickels, dimes, and quarters. If he has five more nickels than dimes and a total of $6, how many of each coin were in his bank?

Found 2 solutions by macston, ankor@dixie-net.com:
Answer by macston(5194)

Insufficient information.

Answer by ankor@dixie-net.com(22740)

let n = no. of nickels
let d = no. of dimes
let q = no. of quarters
:
Write an equation for each statement
:
Dan opened a piggy bank and found a total of thirty-five coins, made up of nickels, dimes, and quarters.
n + d + q = 35
If he has five more nickels than dimes
n = d + 5
and a total of $6,
.05n + .10d + .25q = 6
multiply equation by 4, subtract from the 1st equation
1n + 1d + 1q = 35
.20n + .4d + 1q = 24
----------------------subtraction eliminates q
.80n + .6d + 0 = 11
replace n with (d+5) from the 2nd equation
.8(d+5) +.6d = 11
.8d + 4 + .6d = 11
1.4d = 11 - 4
1.4d = 7
d = 7/1.4
d = 5 dimes
then n = d + 5, therefore
n = 5 + 1
n = 10 nickels
Replace d and n in the 1st equation to find q
10 + 5 + q = 35
q = 35 - 20
q = 20 quarters
:
how many of each coin were in his bank?
10 nickels, 5 dimes, 20 quarters
:
;
:
see if this adds up
10(.05) = .50
5 (.10) = .50
20*.25) = 5.00
----------------
total amt: $6